if two opposite sides of a square are each increased by 3 cm, and the other two sides are each decreased by 2 cm, the area is increased by 8 cm. find the length of a side of the square.
Answer by ankor@dixie-net.com(22740)

if two opposite sides of a square are each increased by 3 cm, and the other two sides are each decreased by 2 cm, the area is increased by 8 cm.
find the length of a side of the square.
:
let s = the side of the original square
then
s^2 = the original area
:
Equation for the new area
(s+3)(s-2) = s^2 + 8
FOIL
s^2 - 2s + 3s - 6 = s^2 + 8
subtract s^2 from both side
-2s + 3s = 8 + 6
s = 14 is the side of the square
;
:
see if that checks out
14^2 = 196 the original area
17 * 12 = 196 + 8
204 = 204